OpenStudy (anonymous):

[6.03] What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 1 3x - 4y = 10 2 -1 1 -2

OpenStudy (anonymous):

2x+3y=1 3x-4y=10 one way to solve this is by multiplying each side of the equation by a number to make either x or y opposite each other...didnt explain that very well so i will show it to you by an equation. 3(2x+3y=1)=6x+9y=3 -2(3x-4y=10)=-6x+8y=-20 then add the two sides together to isolate the variable. 6x+9y=3 -6x+8y=-20 17y=-17 y=-1 then plug that back into the equation 2x+3(-1)=1 2x+-3=1 2x=4 x=2 y=-1 (2,-1)

OpenStudy (anonymous):

Ok, thx so much!